Cola Inc. and Soda Co. are two of the largest and most successful beverage companies in the world in terms of the products that they sell and their receivables management practices. To evaluate their ability to collect on credit sales, consider the following rounded amounts reported in their annual reports (amounts in millions).

Cola Inc. Soda Co.
Fiscal Year Ended: 2015 2014 2013 2015 2014 2013
Net Sales $ 35,519 $ 31,990 $ 32,944 $ 58,838 $ 44,232 $ 44,251
Accounts Receivable 4,498 3,833 3,161 6,487 4,734 3,804
Allowance for Doubtful Accounts 50 57 53 146 92 72
Accounts Receivable, Net of Allowance 4,448 3,776 3,108 6,341 4,642 3,732

Required:

  1. Calculate the receivables turnover ratios and days to collect for Cola Inc. and Soda Co. for 2015 and 2014. (Use 365 days in a year. Do not round intermediate calculations on Accounts Receivable Turnover Ratio. Round your final answers to 1 decimal place. Use final rounded answers from Accounts Receivable Turnover Ratio for Days to Collect ratio calculation.)
Answer #1
COLA INC 2015 2014
ACCOUNT RECEIABLE TURNOVER RATIO=NET SALE/AVERAGE NET ACCOUNT RECEIVABLES 35519/(4448+3776/2)=8.6 31990/(3776+3108/2)=9.3
DAYS TO COLLECT=365/ACCOUNT RECEIAVLE TURNOVER RATIO 365/8.6=42 DAYS 365/9.3=39 DAYS
SODA CO. 2015 2014
ACCOUNT RECEIABLE TURNOVER RATIO=NET SALE/AVERAGE NET ACCOUNT RECEIVABLES 58838/(6341+4642/2)=10.7 44232/(4642+3732/2)=10.6
DAYS TO COLLECT=365/ACCOUNT RECEIAVLE TURNOVER RATIO 365/10.7=34 DAYS 365/10.6=34 DAYS